If π‘Ž = 4𝑑^βˆ’3/2 , 𝑠 = 16 π‘€β„Žπ‘’π‘› 𝑑 = 4, π‘Žπ‘›π‘‘ 𝑠 = 25 π‘€β„Žπ‘’π‘› 𝑑 = 6, find the equation of motion 𝑠 = 𝑓(𝑑) and the velocity function 𝑣(𝑑).


1
Expert's answer
2021-03-01T07:15:40-0500

Acceleration "a" is given as


"a = \dfrac{dv}{dt} \implies dv = a\;dt \implies v = \int a\;dt"

The function of the velocity ("v") is given as:


"v =\int 4t^{-3/2} dt\\
= 4 \int t^{-3/2} dt\\
= 4\int \dfrac{1}{t^\frac{3}{2}} dt\\
= 4 \Bigg(-\dfrac{2}{\sqrt{t}}\Bigg)\\
v=-\dfrac{8}{\sqrt{t}} + c"

Velocity ("v") is given as:


"v = \dfrac{ds}{dt} \implies ds = v\; dt \implies s = \int v\;dt"

The function of the equation of motion can be obtained as follows:


"s= \int \Big(-\dfrac{8}{\sqrt{t}} + c \Big) dt\\
= -8 (2\sqrt{t}) + ct\\
s = - 16\sqrt{t} + ct +c_1"

When s = 16 and t = 4:


"16 = -16(\sqrt{4})+c(4) + c_1\\
= -16(2) + 4c + c_1\\
16= -32 +4c +c_1\\
\implies 4c+c_1 = 48 \qquad\dots (eqn\;1)"

When s = 25 and t = 6:


"25 = -16(\sqrt{6})+c(6) + c_1\\
= -16(\sqrt{6}) + 6c + c_1\\
25= -16(2.45) +6c +c_1\\
6c+c_1 = 25+39.2 \\
\implies 6c+c_1 = 64.2 \qquad\dots (eqn\;2)"

Solving eqn 1 and eqn 2 simultaneously:


"\quad4c+c_1=48\\
(-) 6c+c1=64.2\\
---------\\
-2c=-16.2\\
c= 8.1"

Hence we find "c_1" :


"4(8.1) +c_1 = 48\\
32.4 +c_1 = 48\\
c_1 = 48 -32.4\\
c_1=15.6"

Therefore, the equation of the motion "s = f(t)" is:


"s=f(t)=-16\sqrt{t}+8.1t+15.6"

The velocity "v(t)" is:


"v(t) = -\dfrac{8}{\sqrt{t}} + 8.1"